Please show your work:
$2,350 is invested in account paying 9% compounded semiannually. How much will the account be worth after 8 years? Show your work.

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

$2,350 is invested in account paying 9% compounded semiannually. How much will the account be worth after 8 years?
----------
8 yrs semi-annually = 16 periods
r = 0.09/2 = 0.045
FV = $2350*(1.045)^16
FV = $2350*2.02237
FV = $4752.57

Answer by rfer(16322)

A=P(1+r/n)^nt
A=2350(1+.045)^16
A=2350(2.02237)
A=$4752.57
n=number of compunding periods per year
t=number of years
r=interest rate